OpenStudy (anonymous):

Using the probability distribution table what the probability P(0

OpenStudy (anonymous):

\[\begin{array}{c|c} \text{X}&\text{P}\\ \hline 0&0.2\\ 1&0.4\\2&0.3\\3&0.1 \end{array}\] \[P(0<X<3)=P(1\le X\le2)=P(X=1)+P(X=2)=0.4+0.3\]